OpenStudy (anonymous):

find the derivative of -6/x when x=12..help please...

OpenStudy (amistre64):

you are given a process to work out .. what have you got?

OpenStudy (anonymous):

ok, (-6/(x+h)-(6/x))/h

OpenStudy (amistre64):

youve dropped a negative

OpenStudy (anonymous):

ah so ((-6/(x+h)-(-6/x))/h

OpenStudy (anonymous):

common denom. so:-6x^2-6xh+6x^2+6xh now that's the part I'm confused about

OpenStudy (amistre64):

(-6/(x+h)--6/x)/h (-6x/x(x+h)+ 6(x+h)/x(x+h))/h ( (-6x+6x+6h)/x(x+h) )/h ( (6h)/x(x+h) )/h 6h/hx(x+h) 6/x(x+h)

OpenStudy (amistre64):

\[\lim_0~\dfrac{\dfrac{-6}{x+h}-\dfrac{-6}{x}}{h}\] \[\lim_0~\dfrac{\dfrac{-6x+6(x+h)}{x(x+h)}}{h}\] \[\lim_0~\dfrac{-6x+6(x+h){}}{hx(x+h)}\] \[\lim_0~\dfrac{\cancel{-6x+6x}+6h{}}{hx(x+h)}\] \[\lim_0~\dfrac{6\cancel{h}{}}{\cancel{h}x(x+h)}\to \frac{6}{x(x+0)}\]

OpenStudy (amistre64):

let x=12 :)

OpenStudy (anonymous):

I see it, Thank you very much :DD

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

do you mind helping me on another derivative question?

OpenStudy (anonymous):

f(x)=5x+9 x=2

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

i got (5(x+h)+9)-(5x+9)/h 5x+5h+9-5x-9/h

OpenStudy (amistre64):

the slope of a line is constant ... regardless of any point in the line

OpenStudy (amistre64):

you did fine so far now subtract what gets subtracted

OpenStudy (anonymous):

thank you, so it would be just be 5 then..

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

thank you soo much :DD i think i got this now

OpenStudy (amistre64):

practice helps ;)